# Simulating quantum dynamics in two-dimensional lattices with tensor network influence functional belief propagation

### v2

@@ II.2 Dynamics on tree lattices
It is straightforward to extend the above procedure for finite tree lattices to infinite tree lattices with the same coordination number, known as Bethe lattices [55]. Given the IF-MPS $\mathcal{I}$ for one tree subgraph, the subgraph itself is composed of the tensor $\mathcal{T}$ and $z-1$ IF-MPS $\mathcal{I}$ where $z$ is the coordination number of the Bethe lattice. Fig. 4 illustrates the case when $z=3$. From the above condition, we see that the IF-MPS satisfies a non-linear self-consistent equation.
@@ III Belief propagation with influence functionals
where $\partial j\backslash i$ denotes a set of neighbor sites of $j$ excluding site $i$ and $\cdot$ denotes a tensor contraction. We repeat the propagation until Eq. 1 reaches a fixed point; it forms a local self-consistent BP equation [31, 32],
@@ IV Perfect dephaser point
Figure 7: Tensor network diagram for the dynamics of the kicked Ising model at its perfect dephaser (PD) point. (a) Time evolution is given by a local $X$ rotation (light blue circles) and nearest neighbor $ZZ$ interaction (light orange circles). (b) A dual-unitary property of the self-dual points. White circles indicate a local trace operation. (c) Tensor network diagram of the IF from the light cone construction of the folded TN. The IF is reduced to a product form with the local trace operations recursively, where the IF satisfies the PD property.
@@
We show that these properties hold for Bethe lattices. We prove the PD property from the light cone construction of the folded TN to build the IF in Fig. 7c. The consequent IF has a product form of local trace operations. The proof is based on the recursive structure of the tree. The light-cone IF with time $n$ ($n=4$ in Fig. 7c) has subtrees with time $n-1$. For the recursion, assuming the local trace operation form of IF at time $n-1$, the local trace form of IF at time $n$ is derived from the dual-unitary property (Fig. 7c). A broader class of models also leads to a PD form of the IF if the constituting circuits satisfy the dual-unitary property or a generalization, such as in tri-unitary quantum circuits [62].
@@ V.2.2 Shallow depth circuit results
In Fig. 10, we analyze the convergence behavior of the three methods with respect to bond dimension at $\theta_{h}=7,8,9,10\pi/32$. At $\theta_{h}=7\pi/32$, all three methods converge without the need for bond dimension extrapolation. However, for larger values of $\theta_{h}$, the iPEPS results are not as fully converged at $D=512$ as those obtained from IF-BP and MIX TN, which accounts for the discrepancy observed in Fig. 9. Notably, at $\theta_{h}=9\pi/32$ and $3210\pi/$ 32, the IF-BP result exhibits faster convergence than both iPEPS and MIX TN methods.
@@ VI.1 Method
We now describe a method to include the loop effects of quantum dynamics that are neglected in the IF-BP solution using a cluster expansion, similar to the loop series expansion presented in [82]. We first consider a single cluster region around the local observable of interest and use the BP solution to define its boundary. Fig. 15a illustrates a cluster of maximum loop width $W=2$ surrounded by its BP environment. This cluster idea has previously been used to compute local observable expectation values with PEPS wavefunction [83, 84].
@@ VI.2 Numerical results
We note that the SPD and t-NQG methods are implemented on finite systems. Specifically, SPD is run on an $11\times 1$ 11 lattice with open boundary conditions, while t-NQG uses an $8\times 8$ lattice with periodic boundary conditions. In Fig. 16 at $h=h_{c}$, some differences are observed among the reference data obtained from SPD, iPEPS, and t-NQG. To investigate the origin of these differences, we additionally performed SPD simulations on an $8\times 8$ lattice with periodic boundary conditions, as shown in Fig. 17 up to $t=0.6$. The SPD method contains a controllable threshold parameter $\delta$ that enables us to verify numerical convergence. By decreasing $\delta$ from $2^{-18}$ to $2^{-21}$, we observe that the results on the $8\times 8$ lattice converge to those from the $11\times 1$ 11 system, thereby ruling out finite-size effects as the source of discrepancy. Fig. 17 also includes SPD data with $\Delta t=0.01$ and $\delta=2^{-19}$, taken from Ref. [47]. We note that the SPD method generally requires a smaller $\delta$ to achieve comparable accuracy. Despite this, the SPD data shows good agreement with the iPEPS results up to $t=0.4$, confirming that the visible discrepancies arise from the timestep error. Given the timestep error estimate of $0.003$ from Ref. [47] and the comparison to iPEPS data with $\Delta t=0.01$, most of the remaining deviation of the t-NQG data likely originates from the limited variational expressivity of the particular neural quantum parametrization.
@@
Another factor determining the computational cost of the cluster expansion is the intermediate bond dimension $D_{\text{SU}}$ used in the SU compression and the boundary MPS contraction of the clusters. In our calculations, the boundary MPS bond dimension was set equal to $D_{\text{SU}}$, which provided sufficient accuracy. The required $D_{\text{SU}}$ increases with $t$ to maintain accuracy in the tensor network contraction. Specifically, the required $D_{\text{SU}}$ to converge to the SPD reference data at $h=h_{c}$ is 24, 40, 56, and 64 for $t=0.4,0.6,0.8$, and $1.00.8<t<$ 1.0, respectively. While $D_{\text{SU}}=72$ was used for $t=1.0$, this did not yield fully converged results for large $W$, resulting in an estimated uncertainty of 0.005 in the observable. Consequently, the visible difference between $W=4$ and $W=5$ at $t=1.0$ in Fig. 18 arises not only from loop effects but also from insufficient convergence in $D_{\text{SU}}$.
@@ VII Conclusions
In this work, we introduced the belief propagation (BP) algorithm based on tensor network influence functionals to simulate the dynamics of local observables in quantum lattice systems beyond one dimension. The method is numerically exact on tree lattices in the limit of large bond dimensions and remains accurate on locally tree-like lattices with large loops. Our numerical results validate the effectiveness of IF-BP by reproducing the kicked Ising dynamics quantum experiments on the heavy-hex lattice. Moreover, for longer-time dynamics, the IF-BP approach outperforms traditional tensor network state-based methods, as it leverages the slow, logarithmic growth of temporal entanglement entropy to maintain computational efficiency.

### v3

@@ II.2 Dynamics on tree lattices
It is straightforward to extend the above procedure for finite tree lattices to infinite tree lattices with the same coordination number, known as Bethe lattices [55]. Given the IF-MPS $\mathcal{I}$ for one tree subgraph, the subgraph itself is composed of the tensor $\mathcal{T}$ and z1z-1 IF-MPS $\mathcal{I}$ where $z$ is the coordination number of the Bethe lattice. Fig. 4 illustrates the case when $z=3$. From the above condition, we see that the IF-MPS satisfies a non-linear self-consistent equation.
@@ III Belief propagation with influence functionals
where $\i\partial j\backslash$ i denotes a set of neighbor sites of $j$ excluding site $i$ and $\cdot$ denotes a tensor contraction. We repeat the propagation until Eq. 1 reaches a fixed point; it forms a local self-consistent BP equation [31, 32],
@@ IV Perfect dephaser point
Figure 7: Tensor network diagram for the dynamics of the kicked Ising model at its perfect dephaser (PD) point. (a) Time evolution is given by a local $X$ rotation (light blue circles) and nearest neighbor ZZZZ interaction (light orange circles). (b) A dual-unitary property of the self-dual points. White circles indicate a local trace operation. (c) Tensor network diagram of the IF from the light cone construction of the folded TN. The IF is reduced to a product form with the local trace operations recursively, where the IF satisfies the PD property.
@@
We show that these properties hold for Bethe lattices. We prove the PD property from the light cone construction of the folded TN to build the IF in Fig. 7c. The consequent IF has a product form of local trace operations. The proof is based on the recursive structure of the tree. The light-cone IF with time $n$ ($n=4$ in Fig. 7c) has subtrees with time n1n-1. For the recursion, assuming the local trace operation form of IF at time n1n-1, the local trace form of IF at time $n$ is derived from the dual-unitary property (Fig. 7c). A broader class of models also leads to a PD form of the IF if the constituting circuits satisfy the dual-unitary property or a generalization, such as in tri-unitary quantum circuits [62, 63].
@@ V.2.2 Shallow depth circuit results
In Fig. 10, we analyze the convergence behavior of the three methods with respect to bond dimension at $\theta_{h}=7,8,9,10\pi/32$. At $\theta_{h}=7\pi/32$, all three methods converge without the need for bond dimension extrapolation. However, for larger values of $\theta_{h}$, the iPEPS results are not as fully converged at $D=512$ as those obtained from IF-BP and MIX TN, which accounts for the discrepancy observed in Fig. 9. Notably, at $\theta_{h}=9\pi/32$ and $10\pi/32$, the IF-BP result exhibits faster convergence than both iPEPS and MIX TN methods.
@@ VI.1 Method
We now describe a method to include the loop effects of quantum dynamics that are neglected in the IF-BP solution using a cluster expansion, similar to the loop series expansion presented in [83]. Direct application of [83] to the IF-BP setting is challenging due to the need to construct orthogonal message subspaces, which is complicated by the MPS structure of our messages. Instead, we choose a related approach that is more easily generalized to our setting. A detailed analysis of the cluster expansion approach is done in Ref. [84], focusing on ground-state quantum many-body problems. We first consider a single cluster region around the local observable of interest and use the BP solution to define its boundary. Fig. 15a illustrates a cluster of maximum loop width $W=2$ surrounded by its BP environment. This cluster idea has previously been used to compute local observable expectation values with PEPS wavefunction [85, 86].
@@ VI.2 Numerical results
We note that the SPD and t-NQG methods are implemented on finite systems. Specifically, SPD is run on an $11\times 11$ lattice with open boundary conditions, while t-NQG uses an $8\times 8$ lattice with periodic boundary conditions. In Fig. 16 at $h=h_{c}$, some differences are observed among the reference data obtained from SPD, iPEPS, and t-NQG. To investigate the origin of these differences, we additionally performed SPD simulations on an $8\times 8$ lattice with periodic boundary conditions, as shown in Fig. 17 up to $t=0.6$. The SPD method contains a controllable threshold parameter $\delta$ that enables us to verify numerical convergence. By decreasing $\delta$ from $2^{-18}$ to $2^{-21}$, we observe that the results on the $8\times 8$ lattice converge to those from the $11\times 11$ system, thereby ruling out finite-size effects as the source of discrepancy. Fig. 17 also includes SPD data with $\Delta t=0.01$ and $\delta=2^{-19}$, taken from Ref. [47]. We note that the SPD method generally requires a smaller $\delta$ to achieve comparable accuracy. Despite this, the SPD data shows good agreement with the iPEPS results up to $t=0.4$, confirming that the visible discrepancies arise from the timestep error. Given the timestep error estimate of $0.003$ from Ref. [47] and the comparison to iPEPS data with $\Delta t=0.01$, most of the remaining deviation of the t-NQG data likely originates from the limited variational expressivity of the particular neural quantum parametrization.
@@
Another factor determining the computational cost of the cluster expansion is the intermediate bond dimension $D_{\text{SU}}$ used in the SU compression and the boundary MPS contraction of the clusters. In our calculations, the boundary MPS bond dimension was set equal to $D_{\text{SU}}$, which provided sufficient accuracy. The required $D_{\text{SU}}$ increases with $t$ to maintain accuracy in the tensor network contraction. Specifically, the required $D_{\text{SU}}$ to converge to the SPD reference data within 0.002 at $h=h_{c}$ is 24, 40, 56, and 64 for $t=0.4,0.6,0.8$, and $0.8<t<1.0$, respectively. While $D_{\text{SU}}=72$ was used for $t=1.0$, this did not yield fully converged results for large $W$, resulting in an estimated uncertainty of 0.005 in the observable. Consequently, the visible difference between $W=4$ and $W=5$ at $t=1.0$ in Fig. 18 arises not only from loop effects but also from insufficient convergence in $D_{\text{SU}}$. The required increase in $D_{\text{SU}}$ stems from the entanglement growth within the finite cluster over time, which remains a challenge for the cluster-based classical simulations. We note that, even though the resulting $D_{\text{SU}}$ may appear large from the viewpoint of conventional time-dependent PEPS simulation, in this case it corresponds to the bond dimension in the density operator space, and thus should be compared to the square of the corresponding PEPS bond dimension. In comparison, the maximum bond dimension for iPEPS used in [92] was $D=6$. This employed a corner transfer matrix renormalization group method [96, 97] both to compute the observable expectation values and to optimize the PEPS at each time step, which limits the maximum bond dimension that can be used.
@@ VII Conclusions
In this work, we introduced the belief propagation (BP) algorithm based on tensor network influence functionals to simulate the dynamics of local observables in quantum lattice systems beyond one dimension. The method is numerically exact on tree lattices in the limit of large bond dimensions and remains accurate on locally tree-like lattices with large loops. Our numerical results validate the effectiveness of IF-BP by reproducing the kicked Ising dynamics quantum experiments on the heavy-hex lattice. Moreover, the IF-BP approach outperforms traditional tensor network state-based methods for longer-time dynamics, as it leverages the slow, logarithmic growth of temporal entanglement entropy to maintain computational efficiency. The framework also extends naturally to computing multi-time correlation functions of local observables.